issaashleyg:

Item 3 Classify the number as natural, whole, integer, rational, and/or irrational. Select all terms that are correct. Natural Whole Integer Rational Irrational −25 Natural – −25 Whole – −25 Integer – −25 Rational – −25 Irrational – −25

surjithayer:

\[irrational numbers: \pi,\sqrt{2},\sqrt{3},\sqrt{7}etc.\] \[rational numbers: -5,-2,0,1,2,-15/7,etc \]

surjithayer:

whole nubers:0,1,2,3,... integers: ....,-2,-1,0,1,2,... natural numbers: 1,2,3,....